answersLogoWhite

0

Unfortunately, limitations of the browser used by Answers.com means that we cannot see most symbols. It is therefore impossible to give a proper answer to your question. Please resubmit your question spelling out the symbols as "plus", "minus", "times", "divided by", "equals", "squared", "cubed" etc. There is no symbol between 6h and 9.

User Avatar

Wiki User

∙ 11y ago

What else can I help you with?

Continue Learning about Statistics

If a coin is tossed 5 times what is the probability of getting 5 heads?

If a fair coin is tossed 5 times, the probability of getting 5 heads is:P(H,H,H,H,H) = (1/2)5 = 1/32 = 0.03125 = 3.125% ≈ 3.1%


How many outcomes are there when you flip a coin and roll a die?

12..... h+1, h+2, h+3, h+4, h+5, h+6, t+1, t+2, t+3, t+4, t+5, t+6,


What is probability of throwing 9 heads in a row?

If you toss a coin often enough the probability is 1. The probability of 9 H in the first 9 tosses is (1/2)9 = 1/512


Formula for perimeter of a hexagonal prism?

A hexagonal prism is a polyhedron with two parallel bases bounded by congruent hexagons and with lateral faces bounded by parallelograms that connect corresponding sides of the bases. The height h of the prism is any perpendicular segment drawn from a point on the base to the plane containing the other base. If the segments that join corresponding vertices of the bases are perpendicular to the bases, then the prism is a right hexagonal prism. Otherwise, it is called oblique. The perimeter formula of a right regular hexagonal prism is P = 12s + 6h where s = side and h = height


You are going to toss 2 pennies 20 times how many times do you expect the 2 pennies will come up as 2 heads?

5 times.When tossing 2 pennies there are 4 equiprobable outcomes or events with 1/4 ofprobability for each to happen:P(H,H) = 1/4P(H,T) = 1/4P(T,H) = 1/4P(T,T) = 1/4So if you toss 2 pennies 20 times, the "expected" number of times that 2 heads (H,H)will come up is 1/4 of the 20 times, that is 5 times.

Trending Questions
From a deck of 52 playing cards what is the probability that the card is a queen or a red card? What is the normal d dimer lab value? Find the probability of getting a number greater than 5 when a die is rolled one time? Maria recorded the effects of doing homework on her intelligence. Which of these is a control variable? What is the density of a knife? What term defines chance changes in alelle frequency that have a big effect in small populations? Why are there so many people in the world? What is next in the pattern 33 34 40 41 42 43 44 100 101 102? Can someone provide an explanation on why that is and maybe give an website on where you found it and How come in histograms the bars connect but not in the bar graphs? Can you say the data were obtained? What is 3 over 150 as a percent? What does fnm mean? Two events A and B with probability 0.5 and 0.7 respectively have joint probability of 0.4 The probability that neither A nor B happens is? What does exposure of risk mean? What is the answer for Last airline on which you've flown is it qualitative or quantitative? What is base rate fallacy? What is the approximate percentage score of less than 140 using the 68-95-99.7 rule if a set of test scores is normally distributed with a mean of 100 and a standard deviation of 20? Individuals working in the discipline apply social theories to real life situations? Is Data mining is the process of searching huge amounts of data with the hope of finding a pattern? What is the average weight for a 4 foot 6 inch 10 year old boy?